# Generalized Jeffreys’s approximate objective

### v2

@@ 4 Conclusions
We compared the finite-sample performance of $\mathit{eJAB}$ with Bayes factors computed from MCMC and the Savage–Dickey density ratio for the $t$-test, linear regression, logistic regression, ANOVA, rANOVA, and parametric survival analysis. Under the same generalization of the unit-information prior, $\mathit{eJAB}$ closely approximates the MCMC-based Bayes factor. For a chi-squared test under multinomial or product-multinomial sampling, it agrees with the test-statistic Bayes factor under the alternative, unlike the latter, also registers evidence for the null when true. For nonparametric tests, namely Wilcoxon signed-rank, Mann–Whitney, and Kruskal–Wallis, $\mathit{eJAB}$ performs comparably to Bayes factors based on nonparametric test statistics. Our simulations also show improved performance compared to the BIC approximation to the Bayes factor when $q>1$. Overall, the $\mathit{eJAB}$ calculation is satisfactory, broadly applicable, and very simple.
@@ Acknowledgements
This work was supported by a discovery grant to F.S.N. (RGPIN-04044-2020) from the Natural Sciences and Engineering Research Council of Canada and research accelerator awards to P.V. and Y.L. from the Maud Menten Institute.
@@
Figure 5: Sampling distributions of $D_{n}=-\sqrt{n}\cdot p\ln{p}$ under (A) the local $\mathcal{H}_{1}$ for the one-sample $t$-test and (B) the fixed $\mathcal{H}_{1}$ for the conditional logistic regression.
@@
Figure 7: Comparison of the natural logarithmic BIC approximation to the Bayes factor with alternative Bayes-factor computations at a fixed sample size. The $y$-axis shows the MCMC-based Bayes factor for all panels, except for the two chi-squared tests, which use the test-statistic Bayes factor instead. The red text indicates the effective sample size $n$, except for the Cox model, which shows the number of observations. The dashed lines indicate $\mathit{BF}_{10}$ of 1/3 and 3, representing moderate evidence against and for an effect, respectively.

### v3

@@ 4 Conclusions
We compared the finite-sample performance of $\mathit{eJAB}$ with Bayes factors computed from MCMC and the Savage–Dickey density ratio for the $t$-test, linear regression, logistic regression, ANOVA, rANOVA, and parametric survival analysis. Under the same generalization of the unit-information prior, $\mathit{eJAB}$ closely approximates the MCMC-based Bayes factor. For a chi-squared test under multinomial or product-multinomial sampling, it agrees with the test-statistic Bayes factor under the alternative, unlike the latter, also registers evidence for the null when true. For nonparametric tests, namely Wilcoxon signed-rank, Mann–Whitney, and Kruskal–Wallis, $\mathit{eJAB}$ performs comparably to Bayes factors based on nonparametric test statistics. Our simulations also show improved performance compared to the BIC approximation to the Bayes factor when $q>1$.
@@ Acknowledgements
This work was supported by a discovery grant to F.S.N. (RGPIN-04044-2020) from the Natural Sciences and Engineering Research Council of Canada and research accelerator awards to P.V. and Y.L. from the Maud Menten Institute. P.V. and Z.W. contributed equally to this work. Correspondence concerning this paper should be addressed to F.S.N. (nathoo@uvic.ca).